Peter's Audio has a yield to maturity on its debt of 7.8 percent, a cost of equity of 12.4 percent, and a cost of preferred stock of 8 percent. The firm has 105,000 shares of common stock outstanding at a market price of $22 a share. There are 25,000 shares of preferred stock outstanding at a market price of $45 a share. The bond issue has a total face value of $1.5 million and sells at 98 percent of face value. If the tax rate is 34 percent, what is the weighted average cost of capital?
Answer #1

DETERMINATION OF WEIGHTED AVERAGE COST OF CAPITAL-

Particulars Cost Market Value Weight WACC
Equity 12.4 % $ 22 X 105000=23,10,000 2310000 / 4905000= 0.47 =12.4 X 0.47 = 5.828
Debt 7.8 X 0.66 = 5.148 % $ 1.5 X 10,00,000 X 0.98 = 14,70,000 1470000 / 4905000 = 0.30 = 5.148 X 0.3 = 1.544
Preferred Stock 8 % 25000 X $45 = $ 11,25,000 1125000 / 4905000 = 0.23 8 X 0.23 = 1.84
$ 49,05,000 1.00 9.212 %

THEREFORE WACC = 9.212 %